# Great Models Think Alike and this Undermines AI Oversight

### v1

@@ B.8.1 LM-Judge Prompt without Reference Answer
B.8.2 LM-Judge Prompt with MCQ Options
@@ D. Kidney infections
B.8.4 Open-style CoT Prompt
@@
B.8.5 Coarse Filtering Prompt
@@
B.8.6 Fine-grained Filtering Prompt
@@ C.1 Setup
google/gemma-2-2b (Gemma Team, 2024) anli-r2 (Nie et al., 2019)
@@
microsoft/phi-2 (Li et al., 2023) ethics-utilitarianism (Hendrycks et al., 2020)
@@
wic (Pilehvar & Camacho-Collados, 2018)
@@
Finetuning Methodology: For the above finetuning steps we use Low Rank Adapters (LoRA) (Hu et al., 2021) due to budget constraints, and train a binary classifier the same as Scherlis et al. (2024). We use the confidence weighted loss proposed by Burns et al. (2024). This loss encourages the strong model’s predictions to align with both a weaker model and its own ”hardened” predictions. The hardened predictions are derived by thresholding the strong model’s output. The loss function is defined as:
@@
Similarity vs Weak to Strong Gain: In Figure 13 we reported weak-to-strong gain (Equation 20) on the Y-axis, and similarity ($\kappa_{p}$) on the X-axis. We plot linear grouped by model pair, thus varying the task within each model pair for the linear fit. Figure 13 shows the same scatter points but colored based on the dataset. This shows that weak-to-strong gain is consistently higher for tasks where models are less similar, and how similar two models are depends mostly on the task, i.e. there is not much variance in similarity across the model pairs for a fixed task.

### v2

@@ B.11.1 LM-Judge Prompt without Reference Answer
B.11.2 LM-Judge Prompt with MCQ Options
@@ D. Kidney infections
B.11.4 Open-style CoT Prompt
@@
B.11.5 Coarse Filtering Prompt
@@
B.11.6 Fine-grained Filtering Prompt
@@ C.1 Setup
google/gemma-2-2b (Gemma Team, 2024) anli-r2 (Nie et al., 2020)
@@
microsoft/phi-2 (Li et al., 2023) ethics-utilitarianism (Hendrycks et al., 2021a)
@@
wic (Pilehvar & Camacho-Collados, 2019)
@@
Finetuning Methodology: For the above finetuning steps we use Low Rank Adapters (LoRA) (Hu et al., 2022) due to budget constraints, and train a binary classifier the same way as Scherlis et al. (2024). We use the confidence weighted loss proposed by Burns et al. (2024). This loss encourages the strong model’s predictions to align with both a weaker model and its own “hardened” predictions. The hardened predictions are derived by thresholding the strong model’s output. The loss function is defined as:
@@
Similarity vs Weak to Strong Gain: In Figure 16 we reported weak-to-strong gain (Equation 20) on the Y-axis, and similarity ($\kappa_{p}$) on the X-axis. We group by model pairs, and vary the task within each model pair for the linear fit. Figure 16 shows the same scatter points but colored based on the dataset. This shows that weak-to-strong gain is consistently higher for tasks where models are less similar, and how similar two models are depends mostly on the task, i.e. there is not much variance in similarity across the model pairs for a fixed task.